OpenStudy (anonymous):

If cos theta is less than 0 and cot theta is greater than 0, then the terminal point determined by theta is in: A. Quadrant 2 B. Quadrant 3 C. Quadrant 1 D. Quadrant 4

OpenStudy (jdoe0001):

if cosine is < 0 and cotangent > 0 but \(cot(\theta)=\cfrac{cos(\theta)}{sin(\theta)}\) so if the numerator there is negative, cosine < 0 to get a positive off the fraction, you'd also need a negative sine, that is sine < 0 now on which Quadrant are sine and cosine negative?

OpenStudy (anonymous):

The 3rd quadrant

OpenStudy (jdoe0001):

so, there's the terminal point :)

OpenStudy (anonymous):

Oh I get it. So what if it said cotangent theta > 0. Then what would it be?

OpenStudy (jdoe0001):

well, in this case cotangent is >0, so it'd be the 3rd one :)

OpenStudy (jdoe0001):

now if the cotangent were negative, then what you have would be \(cot(\theta)=\cfrac{-cos(\theta)}{+sin(\theta)}\)

OpenStudy (jdoe0001):

because the fraction, since the cosine is negative, < 0, will need a positive in the denominator to become negative then at Quadrant II cosine is negative and sine is positive

OpenStudy (anonymous):

Thanks soooo much!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):